OpenStudy (anonymous):

a pilot can fly an md-11 aircraft 2160 miles with the wind in the same time as she can fly 1920 miles against the wind is 30, find the speed of the plane in still air

OpenStudy (anonymous):

\[s=d/t\]\[u+30=2160/t\]\[u-30=1920/t\]Solve the simultaneous equations.